The second of two numbers is one more than twice the first. The sum is 25. Find the numbers and show your work.
Answer by ReadingBoosters(3246)

Just in case you are yet in need of the solution...
Set the variables
x = one number
y = other number
...
Given
y = 2x+1
x+y=25
...
Subtitute the first equation into the second
x + 2x+1 = 25
3x +1 =25
Isolate x
3x = 25-1
3x = 24
Divide both sides by 3
x = 8
y = 2(8)+1 = 17
...
The two numbers are 8 and 17
........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.
www.MyHomeworkAnswers.com